Key-Data-Paired Differential Cryptanalysis

Luke Kenneth Casson Leighton

May 22, 2008

Abstract

Traditional Block cipher Differential Cryptanalysis finds relationships between one single bit of a key and a single bit of output, in one fixed location in the output block. The new technique described in this paper, which is being called Key-Data-Paired Differential Cryptanalysis, locates relationships between groups of bits in keys and groups of bits in output blocks (instead of just single bits in fixed locations). Using even as little as 512 blocks of known plaintext on a 128-bit block cipher (for example Rijndael-128), the correlations between groups are sufficient to derive statistical probabilities for individual bits in the key being 0 or 1. Research is underway to ascertain whether those statistical probabilities can be used in a feedback loop to gain better accuracy, by generating weighted keys based on the previous probabilities.

1 1